OpenStudy (whereamiwhere):

Without graphing is the system independent, dependent, or inconsistent? y=2x-8 2x-y=8

OpenStudy (welshfella):

rearrange the second equation to the form y =

OpenStudy (welshfella):

2x - y = 8 Subtract 2x from both sides they divide both sides by -1

OpenStudy (welshfella):

2x - 2x - y = 8 - 2x -y = -2x + 8 now divide both sides by -1

OpenStudy (triciaal):

do you agree this is the same line?